OpenStudy (anonymous):

Write m=4, b=-3 in slope-intercept form.

OpenStudy (anonymous):

The basic form for a slope-intercept equation is. y = mx+b or Y = mx-b if B is a negative integer. Simple enough.

OpenStudy (anonymous):

How do you get x or y from m=4 and b=-3?

OpenStudy (anonymous):

You don't need to get X or Y if you're writing a slope intercept equation. When you have equations that are linear on a graphing plane, the X and Y values are then referred to as their domain and range. And therefore, no singular values would be defined.

OpenStudy (anonymous):

Oh ok then thank you!

OpenStudy (anonymous):

No problem, just remember that these questions are easily done, and if you ever need a domain and range, start plugging in X and Y values. To get a direction of th eline.

OpenStudy (anonymous):

Okay!

OpenStudy (anonymous):

Oh ok then thank you!

OpenStudy (anonymous):

How do you get x or y from m=4 and b=-3?

OpenStudy (anonymous):

I'll help you with that since I have the time. Your final equation would be y = 4x -3 So, to get the x values, plugin in any number for x. Take any integer between -5 and 5. Say,.. y = 4 (2) - 3 y = 5 when x = 2 There you go!